HOT-306 Chief Scientist Daily Reports

Operations completed at Station ALOHA. The Wirewalker, sediment traps and Gas arrays were recovered without any problem. The small boat operation to repair the WHOTS buoy pCO2 equilibrator was successful. The deep sediment traps were recovered without any problems. Winds increased to 7 to 10 kt from the NNW.

Schedule of Activities (HST)

October 13, 2108

2002 - Start S2C14 CTD cast to 1000 m
2107 - End of cast
2120 - Start hand-held net tow
2150 - End net tow
2203 - Start net tow
2231 - End net tow
2300 - Start S2C15 CTD cast to near-bottom

October 14, 2108

0050 - CTD at 7 m off the bottom 22 45.989'N, 157 59.994'W
0236 - End of cast
0251 - Start Optics cast
0423 - End of Optics cast
0430 - Transit to recover Gas array
0600 - Recovering Gas array, 22 37.425'N, 158 2.844'W
0621 - End recovery
0625 - Transit to recover Wirewalker
0645 - Recovering Wirewalker, 22 36.57'N, 158 4.075'W
0700 - End recovery
0705 - Transit to recover sediment traps
0726 - Recovering sediment traps, 22 36.038'N, 158 2.569'W
0758 - End recovery
0800 - Transit to Station 50
0920 - Arrived at Station 50
0935 - Safety briefing for people participating in the small boat operation
0944 - Start small boat operation to repair pCO2 equilibrator on WHOTS-15 buoy
1056 - Operation on WHOTS buoy completed. Equilibrator replaced and tested
1209 - Start hyperpro cast
1226 - Start second hyperpro cast
1244 - Start hand-held net tow
1315 - End net tow
1331 - Start S50C1 CTD yo-yo cast to 200 m, near the WHOTS-15 mooring, 5 cycles completed
1445 - End of cast
1450 - Transit to deep sediment traps mooring location
1630 - Arrived at mooring location. ADCP secured. Start interrogating acoustic releases and release mooring
1740 - Recovering deep sediment traps
1850 - Traps recovered, all equipment on board
1855 - Transit to Pier 35
1858 - ADCP turned on

Fernando S-Mandujano
HOT-306 Chief Scientist
19:30, October 14, 2018

Operations continue on schedule at Station ALOHA. The Gas array was deployed. Winds have been light from the east and later turned from the north, with flat seas.

One of the members of the science party had an accident while preparing the rossete bottles on October 13 at 10:30 AM, he recovered and was released by the captain after an evaluation.

Schedule of Activities (HST)

October 12-13, 2018

1641 - End of cast
1650 - Transit to pump ship's tanks
1830 - Recovering PP array, 22 44.655'N, 158 1.625'W
1852 - End recovery
1909 - Start S2C5 CTD cast to 1000 m
2015 - End of cast
2023 - Start hand-held net tow
2056 - End net tow
2118 - Start S2C6 CTD cast to 1000 m
2224 - End of cast
2240 - Start net tow
2308 - End net tow
2311 - Start net tow
2340 - End net tow
2358 - Start S2C7 CTD cast to 1000 m

October 13, 2018

0057 - End cast
0204 - Start S2C8 CTD cast to 1000 m
0311 - End of cast
0319 - Transit to pump ship's tanks
0431 - Start Gas array deployment
0451 - Gas array deployed, 22 43.7318'N, 158 2.7081'W
0512 - Start S2C9 CTD cast to 1000 m
0620 - End of cast
0634 - Start hand-held net tow
0703 - End net tow
0802 - Start S2C10 CTD cast to 1000 m
0912 - End of cast
0925 - Transit to pump ship's tanks
1030 - Start hand-held net tow
1050 - End net tow
1057 - Start S2C11 CTD cast to 1000 m
1158 - End cast
1211 - Start net tow
1218 - Steaks of Trichodesmium observed around the vessel
1240 - End net tow
1244 - Start net tow
1312 - End net tow
1357 - Start S2C12 CTD cast to 1000 m.
1512 - End of cast
1600 - Start hand-held net tow
1620 - End net tow
1646 - Start S2C13 CTD cast to 1000 m
1800 - End cast
1810 - Transit to pump ship's tanks

Fernando S-Mandujano
HOT-306 Chief Scientist
18:55, October 13, 2018

Operations started at Station ALOHA, and we are currently about 1 hour behind schedule. The wirewalker, sediment traps and primary productivity array were deployed; and the primary productivity array was recovered without any problems. Winds have decreased from 10 to about 6 kt from the ESE, with flat seas.

Schedule of Activities (HST)

October 12, 2018

0026 - Arrived at ALOHA Station.
0056 - Deployed Wirewalker, 22 45.0024'N, 158 2.135'W
0208 - Deployed sediment traps array, 22 44.938'N, 158 1.016'W
0253 - Start S2C1 CTD cast to 200 m
0324 - End of cast
0527 - Deployed Primary Productivity array, 22 46.0426'N, 158 1.489'W.
0604 - Start hand-held net tow
0640 - End net tow
0701 - Start S2C2 CTD deep cast
0848 - Bottom of the cast, 22 44.998'N, 158 0.011'W
1109 - End of cast
1215 - Start net tow
1242 - End net tow
1306 - Start S2C3 CTD cast to 1000 m
1426 - End of cast
1439 - Start hyperpro cast
1511 - End hyperpro
1539 - Start S2C4 CTD cast to 1000 m
1641 - End of cast
1650 - Transit to pump ship's tanks
1830 - Recovering primary productivity array. 22 44.655'N, 158 1.625'W.
1852 - Recovery completed.
1909 - Start S2C4 CTD cast to 1000 m.

Fernando S-Mandujano
HOT-306 Chief Scientist
19:25, October 12, 2018

Operations at Kahe Station were completed slightly behind schedule. Three of the CTD sensors had some problems during the CTD cast at Kahe and have been fixed. We are in transit to Station ALOHA

Schedule of Activities (HST)

October 11, 2108

0922 - All aboard. Depart from Pier 35 with a slight delay due to harbor traffic
1015 - Safety briefing, Science meeting
1050 - Fire and Abandon ship drills
1202 - Arrived at Kahe Station
1218 - Problems with CTD winch and A-frame hydraulics solved before weight cast
1238 - Weight cast to 500 m with 1200 lb weight.
1315 - End of weight cast
1336 - Start hyperpro cast
1408 - End hyperpro cast
1411 - Start hand-held net tow (D. Caron)
1432 - End of net tow
1457 - Start S1C1 CTD cast to 1000 m. The first Oxygen sensor showed large differences as compared the second sensor. Fluorometer trace showed glitches between 300 and 800 m downcast. The transmissometer did not work during the cast, it showed a constant value.
1601 - End of cast
1615 - Transit to ALOHA Station. We expect to arrive between 00:00 and 00:30 on October 12.
1810 - Replaced first CTD oxygen sensor and transmissometer cable. Reseated fluorometer connection. On-board test shows all instruments working.

Fernando S-Mandujano
HOT-306 Chief Scientist
20:30, October 11, 2018